### v2

# Performance of Large Language Models in Numerical vs. Semantic Medical Knowledge: Benchmarking on Evidence-Based Q&As

Abstract
Objective: Clinical problem-solving requires processing of semantic medical knowledge such as illness scripts and numerical medical knowledge of diagnostic tests for evidence-based decision-making. As large language models (LLMs) show promising results in many aspects of language-based clinical practice, their ability to generate non-language evidence-based answers to clinical questions is inherently limited by tokenization. Therefore, we evaluated LLMs’ performance on two question types: numeric (correlating findings) and semantic (differentiating entities) while examining differences within and between LLMs in medical aspects and comparing their performance to humans.
Methods: To generate straightforward multi-choice questions and answers (QAs) based on evidence-based medicine (EBM), we used a comprehensive medical knowledge graph (encompassed data from more than 50,00 peer-reviewed articles) and created the “EBMQA”. EBMQA contains 105,000 QAs labeled with medical and non-medical topics and classified into numerical or semantic questions. We benchmarked this dataset on two state-of-the-art LLMs, Chat-GPT4 and Claude3-Opus. We evaluated the LLMs accuracy on semantic and numerical question types and according to sub-labeled topics. For validation, six medical experts were tested on 100 numerical EBMQA questions.
Results: In an analysis of 24,542 QAs, Claude3 and GPT4 performed better on semantic QAs (68.7% and 68.4%, respectively) than on numerical QAs (63.7% and 56.7%, respectively), with Claude3 outperforming GPT4 in numeric accuracy (p$<$.001). A median accuracy gap of 7% [5-10] was observed between the best and worst sub-labels per topic, with different LLMs excelling in different sub-labels. Furthermore, humans (82.3%) surpassed both Claude3 (64.3%, p$<$.001) and GPT4 (55.8%, p$<$.001) in the validation test.
Conclusions: Both LLMs excelled more in semantic than numerical QAs, with Claude3 surpassing GPT4 in numerical QAs. However, both LLMs showed inter and intra gaps in different medical aspects and remained inferior to humans. Thus, their medical advice should be addressed carefully.
Key words: Large language models, questions and answers, dataset, evidence-based medicine
1Faculty of Medicine, Tel-Aviv University, Tel-Aviv, IL
2Kahun Medical Ltd, Givatayim, Israel
3Faculty of Medicine, Hebrew University of Jerusalem, Jerusalem, IL
4School of Computer Science and Engineering, The Hebrew University of Jerusalem, Jerusalem, IL
5Leumit Health Services, Tel Aviv-Yafo, IL
6The Azrieli Faculty of Medicine, Bar-Ilan University, Safed, IL
7Kaplan medical center, Rehovot, IL
8Brigham and Women’s Hospital, Harvard Medical School, Boston, MA
1 Introduction
Clinical problem-solving requires the processing of data using the clinician’s fund of knowledge in the form of illness scripts [1, 2], most of which is semantic. The statistical weight of relationships between data that define an illness is the numerical equivalent of medical knowledge that is essential for prioritizing diagnostic hypotheses and decision-making [3].
Clinicians develop and use numerical knowledge through original research, and leverage diagnostic support tools for more complex decision-making [4, 5]. However, the explosive amount of medical knowledge and complex healthcare systems are tremendous challenges to high-quality, evidence-based medicine (EBM) [6, 7].
The breakthrough of Large Language Models (LLMs), which process extensive data and encode knowledge from numerous online studies, shows great promise as tools for medical decision support [8, 9]. LLMs provide users with a sense of reliability and accuracy, but evidence shows that they occasionally generate responses that are not based on actual knowledge or give incorrect explanations [10, 11]. In addition, their performance on non-textual knowledge like medical codes is limited [12]. Thus, physicians continue to express skepticism regarding LLMs and their capacity to outperform humans [13].
Several benchmark studies have addressed this subject by focusing on lengthy questions from licensing exams [8, 14] or on datasets derived from medical abstracts that could only be answered with yes/no/maybe [15].
To create a dataset that consists solely of EBM knowledge and is flexible enough to generate both semantic and numeric QA we used the Kahun knowledge graph- a clinically validated artificial intelligence tool that uses a medical, evidence-based knowledge graph. We have developed a methodology to generate QAs from this knowledge graph and created the EBMQA dataset. The dataset comprises 105,000 short multiple-choice questions, based on insights extracted from full-length articles, and aimed to test LLMs ability to assist physicians.
Finally, we benchmarked two state-of-the-art LLMs: Open-AI’s Chat-GPT4 (GPT4) [16], and Antropic’s Claude3 Opus (Claude3) [17], using part of EBMQA. Additionally, we compared their results to medical experts. Thus, we could evaluate the performance of LLMs in both numerical and semantic QA, identify differences within and between LLMs across diverse medical and non-medical domains, and compare their results to humans.
These analyses allowed us to assess whether physicians can trust LLMs.
2 Methods
2.1 EBMQA
2.1.1 Kahun
Kahun is a diagnostic tool based on artificial intelligence and structured knowledge graph technologies. The knowledge graph encompasses more than 50,000 peer-reviewed publications and more than 20,000,000 medical relations that were mapped by medical experts [18]. Kahun’s unique structure and its EBM content serves as a reasonable platform to generate the EBMQA.
2.1.2 Questions Structure
All QAs were derived from Kahun’s knowledge graph. Each question was generated based on data from nodes and edges in the graph and consisted of three main entities: source (usually a disorder that is related to the target), target (usually a symptom or sign that is related to the source), and background (usually a relevant population related to the source). In this study, we refer to source, target, or background as entities. Additionally, the relation between entities (derived from data on the edges) determines the question type and the specific template that was used to generate the question and the answers.
EBMQA is comprised of two types of questions:
  1. 1.
    Numeric QAs - derived from connections between a source and a single target. Those questions deal with choosing the range in which the correct answer resides (Figure 1).
  2. 2.
    Semantics QAs - derived from connections between a source and up to six targets (possible answers). Each QA deals with choosing the most common target/s related to a source given a specific relation (subtype, location, duration, etc) (Figure 1).
Further examples of both numeric and semantics QAs are in Table S1.
Figure 1: The flowchart of the study: From Kahun’s knowledge graph, which references source, target, and background as edges of the graph (1-2), to the EBMQA dataset and the LLM benchmarking (3-4), which includes both numeric and semantic QAs.
2.1.3 Multiple Choice Questions Structure
The questions in the EBMQA are multiple-choice questions. Numeric QAs have one correct answer, while semantic QAs have up to five correct answers. However, if one does not know what the answer to the question is, an “I do not know” (IDK) option was added to all QAs as a possible answer.
2.1.4 Numerical Data and Possible Answers
Each QA is based on numerical data derived from Kahun’s knowledge graph. This includes minimum, maximum, and mid values estimating the connections between medical entities. We employed statistical methods, including median and median absolute deviation (MAD), to categorize answers into meaningful ranges based on their calculated mid-values. Specific methodologies for categorizing these ranges, and detailed statistical backing for each type of QA, are documented in the Appendix.
2.1.5 QA exclusion
EBMQA aims to provide simple short medical QA, therefore, questions related to multiple sources, backgrounds, or targets (except semantic questions) were excluded from the EBMQA. Additionally, QAs that are not related to medical knowledge (such as the average length of a season), were removed. Duplicated questions were excluded from the EBMQA, though the answer for the remaining question is the average of all duplicated Mean values. QAs with all or no correct answers were deleted from the dataset.
2.1.6 Labeling
Each QA in the study was categorized using multiple medical data labels which were derived from standardized medical classifications such as those provided by Snomed CT [19] and Kahun’s medical expertise. These classifications include, but are not limited to, Medical Type, Medical Subject Type, Medical Discipline, and Prevalence. Each QA was also analyzed for its question length and distribution of answers. Details on the specific labeling criteria and categories are described in the Appendix.
2.2 Benchmark Analysis
2.2.1 QA selection and subanalysis
Due to the relatively different structure of semantic QA and its small number of QA, we analyzed all of them separately. Additionally, we subanalyzed this cohort by the number of different correct answers (one to five correct answers). Regarding numeric QA types and in the search for meaningful parameters that might influence LLM’s performance, the benchmark included QAs based on the following medical and non-medical sub-labels. Detailed descriptions of the QA selection criteria, sub-analyses, and the rationale behind each label used in this benchmark are provided in Appendix (Figure S1). Although different in the total number of QA per label, each label contains an identical number of selected QAs per sub-labeled entity. Additionally, each QA was selected once.
2.2.2 LLMs prompting
In this study, we used two state-of-the-art LLMs: Chat-GPT4 (gpt-4-0125-preview), and Claude3-Opus (claude-3-opus-20240229). Both models’ parameters were set to be: temperature= 0 and maximum tokens= 300. All queries were sent to the LLMs using R-studio (version 4.2.2). Further descriptions of the prompts and suitable examples are documented in the Appendix.
2.2.3 Evaluating LLM’s performance
We evaluated LLM’s performance using the following metrics:
  1. Accuracy - the total number of correct answers suggested by the LLM divided by the total answers suggested by the LLM, excluding IDK answers.
  2. Answer rate (AR) - the total number of both correct and wrong answers suggested by the LLM divided by the total answers suggested by the LLM, including IDK answers.
  3. Majority - the distribution of the most frequent correct answers among the total correct answers.
2.2.4 Human and Prompt validation
To validate the benchmark, 100 numerical questions (questionnaire) were randomly selected. Two physicians and four medical students (in their clinical years) answered the questionnaire, first with the IDK possibility and then with guessing the answer to their IDK questions. Their accuracies with and without guessing were compared to LLMs. Additionally, to test both the effects of adding IDK as a possibility and changing the order of answers (including IDK), eight different prompts were tested: four prompts, each with a different order of possible answers, including IDK and four prompts each with different order of possible answers, without IDK.
2.2.5 Analysis and variables
All statistical analyses were performed using R-studio (version 4.2.2). Categorical variables were represented by a percentage while continuous variables were represented by mean and standard deviation (SD) if distrusted normally or by median and interquartile range (IQR) else. The cutoff for statistically significant results and confidence intervals (CI) were set to 95%. Proportions comparison was conducted using the Chi-Square test. For estimating the correlation of two quantitative variables, we employed the Spearman correlation.
2.2.6 Ethics
This study was approved by the Tel Aviv University Ethics Committee (Institutional Review Board (IRB) protocol number 0008527-2).
3 Results
3.1 EBMQA
The EBMQA contains 105,222 QAs. Additionally, each QA pair was labeled according to metadata labels and medical labels.
3.1.1 Medical labels
The EBMQA encompasses diverse medical data types, including a unique count of 7,746 “Disorders”, 2,547 “Signs or Symptoms”, 1,243 “Lab tests”, 885 “Imaging or procedures” data, 474 “Background” data (demographics, habits, family history, etc.), and more (Figure S2A).
Among the Medical subject types, “Disorders” was the most abundant with 45,964 QAs, followed by “Symptoms and Signs” with 30,152 QAs, “Lab test” with 5,966 QAs, and “Imaging or Procedures” with 4,374 QAs. All the other subjects encompass 640 QAs (Figure S2B).
Focusing on Medical Discipline, the EBMQA contains 64,846 relevant QAs: the leading medical discipline was the Digestive system with 9,879 QAs, followed by Cardiovascular system with 7,847 QAs and Infectious diseases with 7,798 QAs. Musculoskeletal system had the least number of QAs- 2,832 (Figure S2C).
Regarding the “Prevalence” label, the median prevalence was 1e-4 and MAD of 9.810102e-05. Among them 36,653 QAs focused on high-prevalence disorders, 22,139 QAs focused on moderate-prevalence disorders, and 2,531 QAs focused on low-prevalence disorders. (Figure S2D)
3.1.2 Metadata labels
EBMQA includes 13 distinct QA types (Figure S2E). The most frequent QA type is “Sensitivity” with 70% (74,140/105,222) of the total QAs, while eight QAs have less than 900 QAs per QA type: Specificity, PLR (positive likelihood ratio), NLR (negative likelihood ratio), Relative Risk, Prevalence, PPV (positive predictive value), NPV (negative predictive value), and Associated Risk.
In total, the median number of words per question (including the question, the instructions and the possible answers) in the EBMQA was 57 [IQR: 53-66] with a MAD of 5. The medium question length group has the majority of QAs, with 59,998 QAs, whereas the short question length group, with 9,968 QAs, represents the minority (Figure S2F). Focusing on each QA type reveals that the “Risk Factor” QA type has the longer median of words per question with 81 [IQR: 80-84] words while the “Sensitivity” QA type has the shorter median words per question with 54 [IQR: 52-58] words (Table S2).
Regarding numeric questions with three range values, the most frequently distributed answer was the mid-range values (46,431 QAs), followed by the low-range values (28,598 QAs) and the high-range values (14,292 QAs) (Figure S3).
3.2 Benchmark Analysis
Out of the 105,222 QA in the EBMQA, the same 24,542 questions were presented to each LLMs. The ”numeric” QA analysis comprised 90% (22,000/24,542) of the questions, whereas ”semantic” QA analysis accounted for the remaining 10% (2,542/24,254).
Both LLMs demonstrated better performances in the semantic QAs than in the numeric QAs in terms of accuracy and AR (Calude3- 68.65% (1,592.78/2,320) vs 61.29 (8,583/14,005), p$<$.001, GPT4- 68.38% (1708.85/2499) vs 56.74% ( 12,038/ 21,215), p$<$.001 and Calude3- 94.62% (2,320/2,542) vs 63.66% (14,005/ 22,000), p$<$.001, GPT4- 98.31% (2,499/2,542) vs 96.4% (21,215/22,000), p$<$.001 respectively). From an inter-model perspective, Calude3 outperformed GPT4 in numeric accuracy though no significant difference was found in semantic accuracy. However, in comparison to Calude3, GPT4 had a higher AR in both semantic and numeric questions (Table1).
Question Type Metric Claude3 GPT4 P-value (Chi-squared) OR CI
Semantic Accuracy 68.65% 68.38% 0.8364 1.01 [0.9-1.14]
(count) (1,592.78/2,320) (1,708.85/2,499)
Answer-rate 94.62% 98.31% $<$.00001 0.18 [0.13-0.25]
(count) (2,320/2,542) (2,499/2,542)
Numeric Accuracy 61.29% 56.74% $<$.00001 1.2 [1.16-1.26]
(count) (8,583/14,005) (12,038/21,215)
Answer-rate 63.66% 96.4% $<$.00001 0.06 [0.06-0.07]
(count) (14,005/22,000) (21,215/22,000)
Table 1: Comparison of Claude3 and GPT4 on Semantic and Numeric Questions
3.2.1 Validation Analysis
Calude3 and GPT4 average accuracy rates, with the IDK option (64.25%,±3.95, 55.75%±1.71, respectively) or without it (59.25%±5, 53.25%±2.89, respectively), were higher than random (33%) or majority guessing (47%) though had lower average accuracy rates in comparing to humans with the IDK option (82.3%±2.82) or without it (78.2%±3.6) (Figure 2, Table S3).
Figure 2: Validation test: Each LLM was tested eight times- four times with the option to “I do not know” (abstain), using the same prompt though in a different order of possible answers, and four times without the abstain option, using the same prompt though in a different order of possible answers. Additionally, six medical experts were tested: first, with the option to abstain, and then without. Confidence intervals of 95% were calculated accordingly while answer-rate (AR) were added only to abstaining instances.
3.2.2 Numeric QA subanalysis
[…]
Figure 3: Numeric QA accuracy and answer-rate sub-labels analysis: (A) Answer distribution, (B) Medical Discipline, (C) Medical Subject type, (D) QA type, (E) Disorders Prevalence, (F) Question length.
4 Discussion
This study aimed to highlight the current gaps in the medical knowledge of LLMs and their current ability to surpass humans. We presented a method to create a QA dataset (EBMQA) from a structured knowledge graph and benchmarked two state-of-the-art LLMs (GPT4 and Claude3) [16, 17]. We demonstrated that both LLMs performed better in semantic QAs than in numerical QAs by asking more than 24,000 QAs (Table1). Claude3 outperformed GPT4 in numerical QAs, and showed similar results in semantic QAs though had significantly lower ARs (Table1). A validation test indicated that the numerical accuracy rates of Claude3 and GPT4 were higher than a majority-guessing, though lower than medical experts (Figure 2).
The use of knowledge graphs for evaluating LLMs is gaining popularity [20, 21, 22]. Kahun’s structured knowledge graph enabled us to generate both semantic and numeric labeled QA, without utilizing advanced models [22]. Additionally, this unique knowledge graph allowed us to create a massive EBM dataset. Moreover, we embraced a data-driven approach in which distractors were based on sub-analysis distribution rather than specific or random values.
The EBMQA, which consists of 105,000 straightforward single-line QAs, was designed to mimic physicians’ strategy of breaking complex medical scenarios into less complicated problems, unlike medical licensing examination datasets which are typically complex-cases oriented [14, 23]. Additionally, the EBMQA addresses numeric/semantic data which is considered fundamental for physicians [24, 25] while dealing with data from articles and embracing the EBM approach [7], as opposed to the abstracted-based yes/no/maybe QAs in PubMedQA [15].
A major concern regarding applying LLMs in healthcare is the uncertainty of providing solid evidence which supports their answers [8]. Clinical evidence predominantly relies on statistical and numerical data. Thus, it is imperative to examine whether LLMs can deliver this type of reasoning. It has been shown that LLMs are more capable when given semantic questions rather than numerical questions though in a relatively low sample size (smaller than 200 QAs) [26]. As far as we know, we were the first to show this trend in the medical field while using a much larger scale (Table1). Furthermore, since both semantic and numeric questions in the EBMQA may address the same entities but from different perspectives, it is questionable whether LLMs can support their semantic answers with statistical data.
While a recent benchmark analysis found that GPT4 outperformed Claude2 in nephrology questions [27], our inter-model examination reveals that Claude3 outperformed GPT4, including in medical disciplines such as neoplastic disorders (Figure 3B). Those results raise the need to constantly benchmark new LLMs as they continuously improve.
Regarding internal model variations, the differences in accuracy between the highest and lowest performing medical disciplines- 8% for Claude3 and 6% for GPT4 (Figure 3B), support previous benchmarks which found that the performance of LLMs can vary across different medical disciplines [27, 28].
Moreover, this comprehensive benchmark widens the medical scope and further supports both intra- and inter-model differences by exploring medical subjects: as Claude3 favors “Lab test” (66%) and struggles with “Disorders” (60%), GPT4 prefers “Imaging and Procedures” (61%) rather than “Lab test” (53%).
As the debate over whether models surpass humans persists [27, 28], the outcomes of our validation tests suggest that humans still excel in certain medical tasks. Therefore, we support further evaluations of LLMs before using them in medical settings.
Furthermore, the insignificant correlation between accuracy and AR, contradicts the theory that a model’s confidence in its response reflects its subject expertise [29]. Thus, abstaining from providing an answer failed to explain the intra-model variance results, specifically in the medical discipline. Therefore, we are concerned that without previous knowledge regarding both the medical discipline and the model, the trustworthiness of LLMs is doubtful. Thus, we encourage further research on this subject.
In terms of prompt engineering, our validation test showed relatively small SDs in prompt accuracy which supported our prompt stability. Moreover, although insignificant, the IDK prompt yielded higher accuracy and, therefore was used.
Our benchmark did not include medical-tuned LLMs, which showed promising results [30], since they are not publicly available. Therefore, a similar benchmark including those LLMs is highly recommended. Moreover, we believe that nowadays, physicians are asking LLMs straightforward questions without adding further information or using external methods like Retrieval-Augmented Generation. Thus, those methods, that might improve the results, are not within the scope of this study, and further research is needed.
5 Conclusions
On the EBMQA dataset, which resembles physicians’ problem-solving approach, LLMs were better at solving semantic than numeric questions. Despite Claude3 surpassing GPT4, both LLMs exhibited inter and intra gaps in medical knowledge. Additionally, humans outperformed both LLMs in numeric questions. These results suggest that LLMs responses, especially numeric ones, should be considered cautiously in clinical settings.
6 Abbreviation
LLM- Large language model ; QA- Question and answer ; IQR- Interquartile range ; MAD- Median absolute deviation ; PLR- Positive likelihood ratio ; NLR- Negative likelihood ratio ; PPV- Positive predictive value ; NPV- Negative predictive value ; EBM- Evidence-based medicine ; IDK- I do not know ; AR- Answer rate ; SD- Standard deviation ; CI- Confidence interval ; IQR- Interquartile range
7 Competing interests
The authors: E.A., M.L., D.H., D.B.J., M.T.K., D.E., S.L., Y.D., S.B., J.M. and S.O. are paid employees by Kahun Ltd. All other authors declare no financial or non-financial competing interests.

### v3

@@
 Methods: To generate straightforward multi-choice questions and answers (QAs) based on evidence-based medicine (EBM), we used a comprehensive medical knowledge graph (encompassed data from more than 50,00 peer-reviewed articles) and created the “EBMQA”.
-EBMQA contains 105,000 QAs labeled with medical and non-medical topics and classified into numerical or semantic questions.
+EBMQA contains 105,222 QAs labeled with medical and non-medical topics and classified into numerical or semantic questions.
 We benchmarked this dataset on two state-of-the-art LLMs, Chat-GPT4 and Claude3-Opus.
@@
 We have developed a methodology to generate QAs from this knowledge graph and created the EBMQA dataset.
-The dataset comprises 105,000 short multiple-choice questions, based on insights extracted from full-length articles, and aimed to test LLMs ability to assist physicians.
+The dataset comprises 105,222 short multiple-choice questions, based on insights extracted from full-length articles, and aimed to test LLMs ability to assist physicians.
 Finally, we benchmarked two state-of-the-art LLMs: Open-AI’s Chat-GPT4 (GPT4) [16], and Antropic’s Claude3 Opus (Claude3) [17], using part of EBMQA.
@@
 In this study, we refer to source, target, or background as entities.
-Additionally, the relation between entities (derived from data on the edges) determines the question type and the specific template that was used to generate the question and the answers.
+Additionally, the relation between entities (derived from data on the edges) determines the question type and the specific template that was used to generate the question and the answers.Further explanation regarding template creation provided in the Appendix.
 EBMQA is comprised of two types of questions:
@@
 Due to the relatively different structure of semantic QA and its small number of QA, we analyzed all of them separately.
-Additionally, we subanalyzed this cohort by the number of different correct answers (one to five correct answers).
-Regarding numeric QA types and in the search for meaningful parameters that might influence LLM’s performance, the benchmark included QAs based on the following medical and non-medical sub-labels.
-Detailed descriptions of the QA selection criteria, sub-analyses, and the rationale behind each label used in this benchmark are provided in Appendix (Figure S1).
-Although different in the total number of QA per label, each label contains an identical number of selected QAs per sub-labeled entity.
-Additionally, each QA was selected once.
+Regarding numeric QA types and in the search for meaningful parameters that might influence LLM’s performance, the benchmark included QAs based on three medical (Medical subject type, Medical Discipline and Prevalence) and three non-medical sub-labels (QA types, Question length and Answers distribution) as further detailed in Appendix (Figure S1).
+All QAs were selected randomly, and although the total number of QAs per label varied, each label contains an identical number of selected QAs per sub-labeled entity.
+Additionally, each QA was selected only once.
 2.2.2 LLMs prompting
@@
 We evaluated LLM’s performance using the following metrics:
-  1. Accuracy - the total number of correct answers suggested by the LLM divided by the total answers suggested by the LLM, excluding IDK answers.
-  2. Answer rate (AR) - the total number of both correct and wrong answers suggested by the LLM divided by the total answers suggested by the LLM, including IDK answers.
+  1. Accuracy - for both semantic and numeric QAs, the total number of correct answers suggested by the LLM divided by the total answers suggested by the LLM, excluding IDK answers.
+  2.
+Answer rate (AR) - for both semantic and numeric QAs, the total number of both correct and wrong answers suggested by the LLM divided by the total answers suggested by the LLM, including IDK answers.
   3. Majority - the distribution of the most frequent correct answers among the total correct answers.
-2.2.4 Human and Prompt validation
-To validate the benchmark, 100 numerical questions (questionnaire) were randomly selected.
-Two physicians and four medical students (in their clinical years) answered the questionnaire, first with the IDK possibility and then with guessing the answer to their IDK questions.
-Their accuracies with and without guessing were compared to LLMs.
-Additionally, to test both the effects of adding IDK as a possibility and changing the order of answers (including IDK), eight different prompts were tested: four prompts, each with a different order of possible answers, including IDK and four prompts each with different order of possible answers, without IDK.
+2.2.4 Prompt sensitivity analysis
+To test both the effects of adding IDK as a possible answer and changing the order of answers (including IDK), eight different prompts were tested on 100 numerical questions that were randomly selected (questionnaire): four prompts, each with a different order of possible answers, including IDK, and four prompts, each with different order of possible answers, without IDK.
+2.2.5 Human validation
+To validate the QAs in the EBMQA, two physicians and four medical students (in their clinical years) answered the questionnaire, first with the IDK possibility and then with guessing the answer to their IDK questions.
+Their accuracies with and without guessing were compared to LLMs’.
 2.2.6 Analysis and variables
-All statistical analyses were performed using R-studio (version 4.2.2).
-Categorical variables were represented by a percentage while continuous variables were represented by mean and standard deviation (SD) if distrusted normally or by median and interquartile range (IQR) else.
-The cutoff for statistically significant results and confidence intervals (CI) were set to 95%.
-Proportions comparison was conducted using the Chi-Square test.
+All statistical analyses were performed using R-studio (R version 4.2.2).
+Categorical variables were represented by a percentage while continuous variables were represented by mean and standard deviation (SD) if distributed normally or by median and interquartile range (IQR) else.
+The cutoff for statistically significant results was set at was set at alpha= 0.05 and confidence intervals (CI) were set to 95%.
+Proportions comparison was conducted using the ”Proportion test”.
 For estimating the correlation of two quantitative variables, we employed the Spearman correlation.
@@
 Table 1: Comparison of Claude3 and GPT4 on Semantic and Numeric Questions
-3.2.1 Validation Analysis
-Calude3 and GPT4 average accuracy rates, with the IDK option (64.25%,±3.95, 55.75%±1.71, respectively) or without it (59.25%±5, 53.25%±2.89, respectively), were higher than random (33%) or majority guessing (47%) though had lower average accuracy rates in comparing to humans with the IDK option (82.3%±2.82) or without it (78.2%±3.6) (Figure 2, Table S3).
+3.2.1 Prompt sensitivity analysis
+[…]
+3.2.2 Human Validation
+Calude3 and GPT4 average accuracy rates, with the IDK option or without it were higher than random (33%) or majority guessing (47%) though had lower average accuracy rates in comparing to humans with the IDK option ($82.3\%\pm 2.82$) or without it ($78.2\%\pm 3.6$) (Figure 2, Table S3).
 Figure 2: Validation test: Each LLM was tested eight times- four times with the option to “I do not know” (abstain), using the same prompt though in a different order of possible answers, and four times without the abstain option, using the same prompt though in a different order of possible answers.
@@
 […]
-Figure 3: Numeric QA accuracy and answer-rate sub-labels analysis: (A) Answer distribution, (B) Medical Discipline, (C) Medical Subject type, (D) QA type, (E) Disorders Prevalence, (F) Question length.
+Figure 3: Numeric QA accuracy and answer-rate sub-labels analysis: (A) Answer distribution, (B) Medical Discipline, (C) Medical Subject type, (D) QA type, (E) Disorders Prevalence, (F) Question length.Red asterisks represent proportion p-values: .05$<$ *$<$.01, ***$<$.0001
 4 Discussion
@@
 Kahun’s structured knowledge graph enabled us to generate both semantic and numeric labeled QA, without utilizing advanced models [22].
-Additionally, this unique knowledge graph allowed us to create a massive EBM dataset.
-Moreover, we embraced a data-driven approach in which distractors were based on sub-analysis distribution rather than specific or random values.
-The EBMQA, which consists of 105,000 straightforward single-line QAs, was designed to mimic physicians’ strategy of breaking complex medical scenarios into less complicated problems, unlike medical licensing examination datasets which are typically complex-cases oriented [14, 23].
+Our QAs generation process, which relies on templates designed to fit a source-target-background graph structure can be applicable to other graphs with a similar structure.
+Additionally, this relatively large knowledge graph allowed us to create a massive EBM dataset.
+Moreover, we embraced a data-driven approach in which distractors were based on sub-analysis distribution rather than specific or random values
+The EBMQA, which consists of 105,222 straightforward single-line QAs, was designed to mimic physicians’ strategy of breaking complex medical scenarios into less complicated problems, unlike medical licensing examination datasets which are typically complex-cases oriented [14, 23].
 Additionally, the EBMQA addresses numeric/semantic data which is considered fundamental for physicians [24, 25] while dealing with data from articles and embracing the EBM approach [7], as opposed to the abstracted-based yes/no/maybe QAs in PubMedQA [15].
@@
 Furthermore, since both semantic and numeric questions in the EBMQA may address the same entities but from different perspectives, it is questionable whether LLMs can support their semantic answers with statistical data.
-While a recent benchmark analysis found that GPT4 outperformed Claude2 in nephrology questions [27], our inter-model examination reveals that Claude3 outperformed GPT4, including in medical disciplines such as neoplastic disorders (Figure 3B).
+A recent benchmark analysis, focused on nephrology QAs only, found that GPT4 outperformed Claude2 [27].
+Although our inter-model examination did not include a direct nephrology compression, due to a different classification method, it reveals that generally Claude3 outperformed GPT4 and specifically in a variety of medical disciplines such as neoplastic disorders, nervous system and more (Figure 3B).
 Those results raise the need to constantly benchmark new LLMs as they continuously improve.
 Regarding internal model variations, the differences in accuracy between the highest and lowest performing medical disciplines- 8% for Claude3 and 6% for GPT4 (Figure 3B), support previous benchmarks which found that the performance of LLMs can vary across different medical disciplines [27, 28].
-Moreover, this comprehensive benchmark widens the medical scope and further supports both intra- and inter-model differences by exploring medical subjects: as Claude3 favors “Lab test” (66%) and struggles with “Disorders” (60%), GPT4 prefers “Imaging and Procedures” (61%) rather than “Lab test” (53%).
+[…]
 As the debate over whether models surpass humans persists [27, 28], the outcomes of our validation tests suggest that humans still excel in certain medical tasks.
@@
 Thus, we encourage further research on this subject.
-In terms of prompt engineering, our validation test showed relatively small SDs in prompt accuracy which supported our prompt stability.
-Moreover, although insignificant, the IDK prompt yielded higher accuracy and, therefore was used.
+In terms of prompt engineering, our sensitivity analysis showed relatively small SDs in prompt accuracy which supported our prompt stability.
+Additionally, although insignificant, the IDK prompt yielded higher accuracy and, therefore was used.
+Moreover, changing the order of the distractors did not significantly affect the performances of the LLMs.
 Our benchmark did not include medical-tuned LLMs, which showed promising results [30], since they are not publicly available.